OpenStudy (anonymous):

what is this x^2-49/x^2+7x in simplified form?

OpenStudy (anonymous):

x + 7 / x - 7

OpenStudy (luigi0210):

\[\frac{ x^2-49 }{ x^2+7x }=\frac{ (x-7)(x+7) }{ x(x+7) }=\frac{ (x-7) }{ x }\]

OpenStudy (anonymous):

so i was wrong??

OpenStudy (anonymous):

@LaShaun97 yea

OpenStudy (anonymous):

that wasnt even an option

OpenStudy (luigi0210):

How did you arrive at that answer @LaShaun97 ?

OpenStudy (anonymous):

@Luigi0210 can u help me with 1 more?

OpenStudy (anonymous):

well actually i had a problem similar to that problem and i just looked at the first couple of numbers and i picked the answer i had, which was the answer to a completely different problem....sorry

OpenStudy (luigi0210):

Sure

OpenStudy (anonymous):

which graph shows the quadratic function y=3x^2-12x+10

OpenStudy (luigi0210):

It is D, I used the derivative to locate a max/min and went from there

OpenStudy (anonymous):

thnx

OpenStudy (luigi0210):

No problem :D